Strassel Investors buys real estate, develops it, and resells it for a profit. A new property is available, and Bud Strassel, the president and owner of Strassel Investors, believes if he purchases and develops this property it can then be sold for $155000. The current property owner has asked for bids and stated that the property will be sold for the highest bid in excess of $100000. Two competitors will be submitting bids for the property. Strassel does not know what the competitors will bid, but he assumes for planning purposes that the amount bid by each competitor will be uniformly distributed between $100000 and $145000. Develop a worksheet that can be used to simulate the bids made by the two competitors. Strassel is considering a bid of $125000 for the property. Using a simulation of 1000 trials, what is the estimate of the probability Strassel will be able to obtain the property using a bid of $125000? Round your answer to 1 decimal place. Enter your answer as a percent. %

How much does Strassel need to bid to be assured of obtaining the property? $ 145,000

What is the profit associated with this bid?

Use the simulation model to compute the profit for each trial of the simulation run. With maximization of profit as Strassel’s objective, use simulation to evaluate Strassel’s bid alternatives of $125000, $135000, or $145000. What is the recommended bid, and what is the expected profit? A bid of results in the largest mean profit of $ .
Answer #1

The simulation model is following:

EXCEL FORMULAS:

Cell Formula Copy to
B11 =$C$6+RAND()*($C$7-$C$6) B11:C1010
D11 =MAX(B11:C11) D11:D1010
E11 =IF(D11<$F$3,1,0) E11:E1010
F11 =($C$3-$F$3)*E11 F11:F1010
F5 =AVERAGE(E11:E1010)
F7 =AVERAGE(F11:F1010)

1)

Estimate of probability of obtaining the property using a bid of $ 125,000 = 29.8 %

Profit associated with bid value of $ 125,000 = $ 8,940

2)

The upper limit of competitors bid is $ 145,000.

Therefore, Strassel need to bid $ 145,000 to be 100% sure of obtaining the property.

3) Profit associated with this bid = 155000 - 145000 = $ 10,000

4)

Expected profit with bid of $ 125,000 = $ 8,940 (as determined in part 1)

Expected profit with bid of $ 135,000 = $ 12,320

Expected profit with bid of $ 145,000 = $ 155000 - 145000 = 10,000

From the above simulation results, it is clear that the highest profit is achieved with a bid value fo $ 135,000

ANSWER $ 135,000

Largest mean profit = $ 12,320